Chargaffs results yielded a molar ratio of 1.45 for A to G in hen DNA, 1.29 for T to C, 1.06 for A to T, and 0.910 for G to C. The ratio of purines to pyrimidines is 0.990. Calculate the mole fractions of A, G, T and C in hen DNA Mole fraction of A Mole fraction of G Mole fraction of T : Mole fraction of C
Answer #1

Given that Purines: pyrimidine is 0.990. The other given ratios are as follows: A:G=1.45 T:C=1.29 A:T=1.09 G:C=0.910 So, A+G=

As known, A+G+C+T=1....(2) A:T=1.09 A = 1.091 Substitute the values of A, C and C in terms of T in equation (2), A+G+C+T=1 1.C=0.771 C=0.77 (0.287) = 0.221 Therefore, the mole fractions are as follows: A = 0.313 G=0.178 C=0.221 T= 0.287